# ReLI: A Language-Agnostic Approach to Human-Robot Interaction

### v1

@@ III-C Command Decision and Parsing
Additionally, to ensure that the interpreted action plan accurately reflects the user’s intent, especially for complex long-horizon tasks, we introduce a security confirmation mechanism that prompts the user to approve or reject the generated action plans before being deployed for execution. We modelled the confirmation as a binary decision problem, $\rho_{d}\in\{0,1\}$, inferred by applying a learned linear classifier $\mathbf{v}$ to an embedding $\psi(r)$ of the LLM’s interpretation $r$, as shown in Eq. (5):
@@ III-D Visuo-lingual Perception and Object Localization
The ReLI’s visuo-lingual pipeline (bottom of Fig. 2) relies on foundation vision-language models, e.g., CLIP [74] and zero-shot segmentation models (SAM [37]), augmented with geometric depth fusion and uncertainty-aware classification to ground linguistic references into spatially localised entities in the robot’s operational environment. Formally, let $\mathcal{V}_{s}=\{(\mathcal{I}_{t},\mathcal{D}_{t},u_{t})\}_{t=1}^{T}$ be the sequence of time-synchronized RGB-D frames and odometry signals from the robot’s observation sensors, where $\mathcal{I}_{t}\in\mathbb{R}^{H\times W\times 3}$ is the stream of RGB frames, $\mathcal{D}_{t}\in\mathbb{R}^{H\times W}$ is the corresponding depth map, and $u_{t}$ encodes the transformations in the robot’s local frame at time $t$. For each $\mathcal{I}_{t}$, we employ SAM [37] to generate $N$ candidate masks $\{\mathcal{M}_{i}\}_{i}^{N}$ through prompt-guided and automatic segmentation.
@@
In Eq.(6), a higher $\tau$ increases the model’s confidence, whereas a lower $\tau$ yields a smoother distribution, with greater uncertainty. To ensure that only confident predictions propagate downstream, we filter uncertain detections through an energy-based uncertainty quantification score, $\mathcal{E}(\mathcal{M}_{i})=-\text{log}\sum_{j}\exp\!\bigl{(}S_{ij}\bigr{)}>\tau_{e}$ by rejecting masks exceeding the defined energy threshold $\tau_{e}$. In reality, the environment’s perception quality depends heavily on the accuracy of the observation sources. To account for potential degradations of the robot’s task environment (e.g., low light, occlusions, motion blur), we introduce a function $\Theta_{ij}(\Omega)$ that incorporates environmental conditions for degradation-aware weighting. We downweight probabilities for masks in degraded regions using $\Theta_{ij}(\Omega)=\exp\bigl{(}-\beta\,\|\Omega_{i}\|\bigr{)}$, where $\|\Omega_{i}\|$ quantifies local adverse conditions, and $\beta$ regulates the sensitivity. Formally, Eq. (6) can be rewritten as depicted in Eq. (7):
@@
where $\hat{\mathcal{D}}_{\text{mono}}$ is the MiDaS [7] monocular depth prediction. Therefore, for the detected object $o_{j}$ with mask centroid $(u_{c}\,,v_{c})$ at depth $z_{c}$, we apply a pinhole camera model to back-project the pixel into 3D space, $\mathbf{x}_{o}\in\mathbb{R}^{3}$, i.e., $\mathbf{x}_{o}=\Pi^{-1}(u_{c}\,,v_{c},z_{c}).$ We then transform $\mathbf{x}_{o}$ to the robot’s base frame using iterative TF lookups to handle temporal synchronisation. Simultaneously, we used the Kalman filter to track the object poses, modelling the state dynamics as $\mathbf{x}_{t+1}=\mathbf{F}\mathbf{x}_{t}+\mathbf{w}$ to smooth pose estimates and account for motion uncertainty, where $\mathbf{F}$ is the motion model.
@@ III-E Action Execution and Control
We translate the high-level intents obtained from the DnPP (Subsection III-C) into physical robot actions through the action execution mechanism (AEM), see Fig. 2, top right. Generally, the AEM manages all the navigation tasks as well as coordinates the rest of the fundamental robot manoeuvres, sensor-based commands, and safety measures. When the user command requires navigation to coordinates $(x_{g},y_{g},z_{g})$, e.g., “go to the locations (6, 0, 0) and (2, 2, 0)” or to a pre-defined goal destinations, e.g., “go to the Secretary’s office”, we rely on a higher-level motion planning stack to accomplish these tasks. First, we employ a highly efficient Rao-Blackwellized particle filter-based algorithm [17] to learn occupancy grid maps from the robot’s operational environment. For details on these probabilistic simultaneous localisation and mapping (SLAM) methods, we refer the reader to [90].
@@ V-A Quantitative Results
In terms of specific language observations, ReLI handled instructions in English, Spanish, and a few others nearly perfectly, with average IPA $>$ 99%. We attribute this high performance primarily to their large training corpora and well-established linguistic resources, which enhanced the pre-trained LLM prediction accuracy and action parsing. Conversely, certain languages, e.g., Arabic, Chinese, etc, lagged slightly behind other Indo-European high-resource languages. This discrepancy is attributed to the complexities associated with inputting the logographic characters in our interaction interface. Obviously, other than the vocal or audio commands, we relied completely on the translated task instructions for such languages. Furthermore, the TSR for English and Spanish remained consistent with its highest IPA. French, German, etc, also remained slightly above 97% accuracy. Most Indo-European languages, along with Chinese, Japanese, Swahili, Malay, etc., maintained average response times between $2.10-2.20$ seconds, which is ideally a rapid response time for a multilingual system.

### v2

@@ III-C Command Decision and Parsing
Additionally, to ensure that the interpreted action plans accurately reflects the user’s intent, especially for complex long-horizon tasks, we introduce a security confirmation mechanism that prompts the user to approve or reject the generated action plans before being deployed for execution. We modelled the confirmation as a binary decision problem, $\rho_{d}\in\{0,1\}$, inferred by applying a learned linear classifier $\mathbf{v}$ to an embedding $\psi(r)$ of the LLM’s interpretation $r$, as shown in Eq. (5):
@@ III-D Visuo-lingual Perception and Object Localization
The ReLI’s visuo-lingual pipeline (bottom of Fig. 2) relies on foundation vision-language models, e.g., CLIP [73] and zero-shot segmentation models (SAM [38]), augmented with geometric depth fusion and uncertainty-aware classification to ground linguistic references into spatially localised entities in the robot’s operational environment. Formally, let $\mathcal{V}_{s}=\{(\mathcal{I}_{t},\mathcal{D}_{t},u_{t})\}_{t=1}^{T}$ be the sequence of time-synchronized RGB-D frames and odometry signals from the robot’s observation sensors, where $\mathcal{I}_{t}\in\mathbb{R}^{H\times W\times 3}$ is the stream of RGB frames, $\mathcal{D}_{t}\in\mathbb{R}^{H\times W}$ is the corresponding depth map, and $u_{t}$ encodes the transformations in the robot’s local frame at time $t$. For each $\mathcal{I}_{t}$, we employ SAM [38] to generate $N$ candidate masks $\{\mathcal{M}_{i}\}_{i=1}^{N}$ through prompt-guided and automatic segmentation.
@@
From Eq.(6), a higher $\tau$ increases the model’s confidence, whereas a lower $\tau$ yields a smoother distribution, with greater uncertainty. To ensure that only confident predictions propagate downstream, we filter uncertain detections through an energy-based uncertainty quantification score, $\mathcal{E}(\mathcal{M}_{i})=-\text{log}\sum_{j}\exp\!\bigl{(}S_{ij}\bigr{)}>\tau_{e}$ by rejecting masks exceeding the defined energy threshold $\tau_{e}$. In reality, the environment’s perception quality depends heavily on the accuracy of the observation sources. To account for potential degradations of the robot’s task environment (e.g., low light, occlusions, motion blur), we introduce a function $\Theta_{ij}(\Omega)$ that incorporates environmental conditions for degradation-aware weighting. We downweight probabilities for masks in degraded regions using $\Theta_{ij}(\Omega)=\exp\bigl{(}-\beta\,\|\Omega_{i}\|\bigr{)}$, where $\|\Omega_{i}\|$ quantifies local adverse conditions, and $\beta$ regulates the sensitivity. Formally, Eq. (6) can be rewritten as depicted in Eq. (7):
@@
where $\hat{\mathcal{D}}_{\text{mono}}$ is the MiDaS [7] monocular depth prediction. Therefore, for the detected object $o_{j}$ with mask centroid $(u_{c}\,,v_{c})$ at depth $z_{c}$, we apply a pinhole camera model to back-project the pixel into 3D space, $\mathbf{x}_{o}\in\mathbb{R}^{3}$, i.e., $\mathbf{x}_{o}=\Pi^{-1}(u_{c}\,,v_{c},z_{c}).$ We then transform $\mathbf{x}_{o}$ to the robot’s base frame using iterative TF lookups to handle temporal synchronisation. Simultaneously, we used the Kalman filter to track the object poses, modelling the state dynamics as $\mathcal{X}_{t+1}=\mathrm{F}\mathcal{X}_{t}+\mathrm{w}$ to smooth pose estimates and account for motion uncertainty, where $\mathrm{F}$ is the motion model.
@@ III-E Action Execution and Control
We translate the high-level intents obtained from the DnPP (Subsection III-C) into physical robot actions through the action execution mechanism (AEM), see Fig. 2, top right. Generally, the AEM manages all the navigation tasks and coordinates the rest of the fundamental robot manoeuvres, sensor-based commands, and safety measures. When the user command requires navigation to coordinates $(x_{g},y_{g},z_{g})$, e.g., “go to the locations (6, 0, 0) and (2, 2, 0)” or to pre-defined goal destinations, e.g., “go to the Secretary’s office”, we rely on a higher-level motion planning stack to accomplish these tasks. First, we employ a highly efficient Rao-Blackwellized particle filter-based algorithm [18] to learn occupancy grid maps from the robot’s operational environment. For details on these probabilistic simultaneous localisation and mapping (SLAM) methods, we refer the reader to [90].
@@ V-A Quantitative Results
In terms of specific language observations, ReLI handled instructions in English, Spanish, and a few others nearly perfectly, with an average IPA $>$ 99%. We attribute this high performance primarily to their large training corpora and well-established linguistic resources, which enhanced the pre-trained LLM prediction accuracy and action parsing. Conversely, certain languages, e.g., Arabic, Chinese, etc, lagged slightly behind other Indo-European high-resource languages. This discrepancy is attributed to the complexities associated with inputting the logographic characters in our interaction interface. Obviously, other than the vocal or audio commands, we relied completely on the translated task instructions for such languages. Furthermore, the TSR for English and Spanish remained consistent with its highest IPA. French, German, etc, also remained slightly above 97% accuracy. Most Indo-European languages, along with Chinese, Japanese, Swahili, Malay, etc., maintained average response times between $2.10-2.20$ seconds, which is ideally a rapid response time for a multilingual system.